OpenStudy (aabomosalam1998):

For a polynomial p(x), the value of p(3) is -2. Which of the following must be true about p(x)?

OpenStudy (aabomosalam1998):

A) x-5 is a factor of p(x) B) x-2 is a factor of p(x) C) x+2 is a factor of p(x) D) The remainder when p(x) is divided by x-3 is -2.

jimthompson5910 (jim_thompson5910):

Remainder Theorem: if p(x) is divided by x-k, and the remainder is r, then p(k) = r For example, if p(10) = 9 then p(x) divided by x-10 leads to a remainder of 9

jimthompson5910 (jim_thompson5910):

If the remainder is 0, then x-k is a factor of p(x)

OpenStudy (aabomosalam1998):

Not enough information is provided on whether x=5, x=2 or x=-2 are roots of p(x). I am left with D but I don't fully get why is it true. P(x)/x-k=q(x)+r Where q(x) is the quotient and r is the remainder, but can't the remainder be a polynomial as well.

jimthompson5910 (jim_thompson5910):

`but can't the remainder be a polynomial as well.` any single number without variables attached is technically a polynomial. It's a polynomial of degree 0. So the remainder is always a polynomial

jimthompson5910 (jim_thompson5910):

p(x)/(x-k) = q(x) + r/(x-k) If you multiply both sides by (x-k) and simplify, you'll get p(x) = (x-k)*q(x) + r

jimthompson5910 (jim_thompson5910):

now plug in x = k p(x) = (x-k)*q(x) + r p(k) = (k-k)*q(k) + r ... replace every x with k p(k) = (0)*q(k) + r p(k) = 0 + r p(k) = r

OpenStudy (aabomosalam1998):

What I have meant was that r could be a polynomial of a degree of more than zero. Why is it r/x-k?

jimthompson5910 (jim_thompson5910):

it's r/(x-k) to indicate a fractional portion. For instance 10/7 = 1 + 3/7 the '1' is the quotient q(x) the '3' is the remainder r(x)

jimthompson5910 (jim_thompson5910):

we can multiply both sides of `10/7 = 1 + 3/7` by 7 to get 10/7 = 1 + 3/7 7*(10/7) = 7*(1 + 3/7) 10 = 7*(1) + 7*(3/7) 10 = 7*(1) + 3 The "7" plays the role as the "x-k" The "1" is the quotient The "3" is the remainder

jimthompson5910 (jim_thompson5910):

the remainder r could be some higher degree polynomial, but only if you're dividing polynomials like (x^3+6x+5) over (x-3) notice we have a cubic over a first degree polynomial. The degree of the remainder could be as high as 2

OpenStudy (aabomosalam1998):

Got it, thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem